In each of Problems 1 through 4 draw a direction field for the given differential equations. Based on the direction field, determine the behavior of y as t → +∞. If this behavior depends on the initial value of y at t = 0, describe this dependency.

1. y ' = 3 + 2y

2. y ' = 3 − 2y

3. y ' = −y(5 − y)

4. y ' = y(y − 2)2
Answer #1

④ y a 3+27 y-29=3 this is the first-order ode. . If= ě 2* gert eztakte Yezz - 2 / 2 + cet as tas too y(t) ② Y= 3-24 I J+2y=3Q.1 st0.2 5Q.3 5 -Q.4 -5